OpenStudy (anonymous):

Find the lenght of time for: $500 to be intrest on $1800 at 6%p.a Find the rate percent per annum for: 15000 to be interest on $5400 for 5 years Find the principal required for The simple intrest to be $900 on an amount invested for 2 years at 10% p.a.

OpenStudy (kropot72):

Assuming simple interest, the interest after 1 year is 1800 * 0.06. The number of years to reach a total amount of $500 interest is given by: \[\frac{500}{1800\times0.06}=you\ can\ calculate\]

OpenStudy (anonymous):

So the answer is 4.63 right or125/27

OpenStudy (kropot72):

Correct!

OpenStudy (anonymous):

Thanks do uo you know how to do the others

OpenStudy (kropot72):

Please check the values of the interest and the principal in the second part.

OpenStudy (kropot72):

In the third part, let the principal amount be P. 900 = P * 0.1 * 2 \[P=\frac{900}{0.1\times2}=you\ can\ calculate\]

OpenStudy (anonymous):

1500 to be interest on $5400 for 5 years

OpenStudy (anonymous):

sorry got the Question wrong for second part

OpenStudy (kropot72):

For the second part, assuming simple interest, let the interest rate as a decimal be r. 1500 = 5400 * r * 5 \[r=\frac{1500}{5400\times5}=you\ can\ calculate\]

OpenStudy (anonymous):

=6%

OpenStudy (kropot72):

For the second part, I get a yearly interest rate of approximately 5.55%.

OpenStudy (anonymous):

Hey i got another question $3200 at 10.25% p.a for 2 1/2 years compounded - 3 monthly

OpenStudy (kropot72):

\[A=P(1+\frac{r}{n})^{nt}\] \[A=3200(1+\frac{0.1025}{4})^{2.5\times4}\]

OpenStudy (anonymous):

Wait how did u get 4

OpenStudy (kropot72):

In the general equation, n is the number of compounding intervals per year. In your question the interest is compounded 3-monthly, giving 4 compounding intervals per year.

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

so the answer is 1053655670

OpenStudy (anonymous):

ops wrong answer